# The Covariant Stone-von Neumann Theorem for Locally Compact Quantum Groups

### v1

@@ 3. Heisenberg Representations
Modular representations for group actions on operator algebras already appear in [HI] and [modular], emphasizing pointwise notions of covariant representations. Representations of coactions lack such a pointwise notions, and require a distinct but parallel theory, which is more appropriate from the lens of locally compact quantum groups. Below, we offer a definition of Heisenberg representations which extends the acceptable dynamics and unifies the classical setting under the common theory of corepresentations.
@@
Let $(\mathbb{G},A,\alpha)$ be a left coaction. A Heisenberg representation $(X,\rho^{c},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$ consists of a Hilbert $A$-module $X$ and non-degenerate ***-homomorphisms
@@
be non-degenerate ***-homomorphisms. The following are equivalent:
@@
    $(\widehat{\rho},\rho^{c})$ is a covariant representation of the right coaction $(C^{u}_{0}(\widehat{\mathbb{G}}),\widehat{\mathbb{G}},({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u})$ on $X$.
@@
be non-degenerate ***-homomorphisms. The following are equivalent:
@@
    $(\pi,\rho^{c})$ is a covariant representation of the right coaction $(A,\widehat{\mathbb{G}},{\rm triv})$ on $X$.
Here, ${\rm triv}\colon A\rightarrow M(A\otimes C_{0}(\widehat{\mathbb{G}}))$ denotes the right trivial coaction given by ${\rm triv}(a)=a\otimes 1$.
@@
Consider a Heisenberg $(\mathbb{G},A,\alpha)$-representation $(X,\rho^{c},\hat{\rho},\pi)$ comprising $A=\mathbb{C}$, $X=H$ a Hilbert space, and $\mathbb{G}=G$ a classical abelian locally compact group. Then $\rho^{c}$ and $\widehat{\rho}$ correspond to unitary representations $\widehat{U}$ and $U$ of $\widehat{G}$ and $G$ on $H$ respectively. The conditions (R1) and (R3) are automatically satisfied. Let ${\rm ev}_{g}\colon C_{0}(G)\to\mathbb{C}$ and ${\rm ev}_{\chi}\colon C_{0}(\widehat{G})\to\mathbb{C}$ be the evaluation functionals at points $g\in G$ and $\chi\in\widehat{G}$ respectively. After applying $({\rm ev}_{g}\otimes{\rm id}\otimes{\rm ev}_{\chi})$ to the defining equation of condition (R2), we obtain
@@
Rieffel refined and generalized Mackey’s pioneering work on induced representations of locally compact groups to the setting of of $C^{*}$-representation theory (see the introduction to [rinduced] for a nice history), which recovers the Mackey-Stone-von Neumann Theorem treated in Example 3.4 as a special case of Mackey’s Imprimitivity Theorem (follow the references in [Rieffel]*Paragraph Three). In the latter article, Rieffel defines Heisenberg $G$-modules for an arbitrary locally compact group $G$. These are exactly our Heisenberg $(\mathbb{G},A,\alpha)$-representations for $A=\mathbb{C}$ and $\mathbb{G}$ the dual to a classical locally compact group. The discerning will notice that the integrand in [Rieffel]*Equation (3) has striking similarities to the application of the left regular corepresentation interpreted in the classical setting, an antecedent to our current work.
@@
In [mrw], the authors introduced the notion of a Heisenberg pair which Roy and Timmermann later generalized in [RT18]. Meyer, Roy, and Woronowicz define Heisenberg pairs in the context of locally compact quantum groups derived from manageable multiplicative unitaries (see, e.g. [Wor96]) and are closely related to our Heisenberg representations in the sense of Kustermans and Vaes. Compare for instance the parallels between our Proposition 3.9 and [RT18, Theorem 6.2]. It is known that Kustermans and Vaes’s description of locally compact quantum groups subsume the Woronowicz framework.
We end this section with a theorem which relates our Heisenberg modular representations to dynamics of the crossed product. Much as full crossed products completely capture the dynamics of the associated system, our Heisenberg representations retain the data involved in crossed product constructions.
@@
Let $(\mathbb{G},A,\alpha)$ be a left coaction and $X$ a Hilbert $A$-module. There is a one-to-one correspondence between Heisenberg representations of $(\mathbb{G},A,\alpha)$ on $X$ and covariant representations of the left coaction $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$ on $X$.
@@
$=\chi((\widehat{\rho}\otimes{\rm id})({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u}(h))=\chi((\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\widehat{\rho}(h)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*})$
@@
4. The Covariant Stone-von Neumann Theorem
@@
Let $(\mathbb{G},A,\alpha)$ be a left coaction of $\mathbb{G}$ on $A$. A Schrödinger representation is a quadruple $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha)$, where
@@
are the canonical surjective ***-homomorphisms.
@@
$=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}(\mathrm{V}^{\widehat{\mathbb{G}}}_{23})^{*}\mathrm{V}^{\widehat{\mathbb{G}}}_{23}=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}=(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}\,.$
@@
Using the machinery we have gathered so far, we can prove the main result of our work.
@@
Let $\left(\mathbb{G},A,\alpha\right)$ be a left coaction, where $A$ is an elementary $C^{*}$-algebra, $\mathbb{G}$ is a regular locally compact quantum group, and $\alpha$ is a maximal coaction. Then every Heisenberg $\left(\mathbb{G},A,\alpha\right)$-representation is a multiple of the Schrödinger $\left(\mathbb{G},A,\alpha\right)$-representation.
@@
Consider a Heisenberg $(\mathbb{G},A,\alpha)$-representation $(X,\rho^{\textnormal{c}},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$. Then Proposition 3.9 yields a covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ for $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$. In turn, the covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ determines a nondegenerate representation of the bicrossed product
@@
Since $\alpha$ is maximal, we have an isomorphism $(\lambda^{c}\otimes 1)\ltimes(\hat{\lambda}\otimes 1)\ltimes\alpha\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\to\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$. Abbreviating the isomorphism by $\Phi$, we witness $X$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence with left action coming from the ***-homomorphism $(\rho^{c}\ltimes\hat{\rho}\ltimes\pi)\circ\Phi^{-1}$.
Applying this analysis to the Schrödinger representation $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha),$ we can also see that the correspondence $L^{2}(\mathbb{G})\otimes A$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence. It is in fact an imprimitivity bimodule. To see this, notice that the map $\Phi$ in the previous paragraph agrees with the bicrossed product ***-homomorphism associated to the Schrödinger representation, so that the left action of $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ is given by the identity. Since $A$ and $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ are both simple, the fullness of $L^{2}(\mathbb{G})\otimes A$ is clear, and for the compatibility of the inner products, we compute for elementary tensors
@@
Thus we may apply Proposition 2.1 to deduce that $X$ is a multiple of $L^{2}(\mathbb{G})\otimes A$. We suppress the unitary which implements the equivalence $X\cong\bigoplus_{S}L^{2}(\mathbb{G})\otimes A$, and proceed by identifying these two modules.
@@
recovering the covariant representation $(\rho^{\textnormal{c}},\hat{\rho}\ltimes\pi)$ of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\hat{\alpha}^{\textnormal{op}}).$ Applying Proposition 3.9, we recover the Heisenberg $(\mathbb{G},A,\alpha)$-representation $(\oplus_{S}\big{(}L^{2}(\mathbb{G})\otimes A\big{)},\rho^{\textnormal{c}},\hat{\rho},\pi).$ Since each summand is invariant under the ***-representations, this concludes the proof. ∎
In light of Theorem 4.3, we introduce a new property for a left coaction $(\mathbb{G},A,\alpha)$, which is a quantum-dynamical analog of the behavior described in [Rieffel]*Theorem and [wcrossed]*Theorem 4.29.
@@
A dynamical system $(\mathbb{G},A,\alpha)$ is called Stone-von Neumann unique provided that every Heisenberg representation is a multiple of the Schrödinger representation.
By Theorem 4.3 if $\mathbb{G}$ is regular, $A$ is elementary and $\alpha$ is maximal, then $(\mathbb{G},A,\alpha)$ is Stone-von Neumann unique. In the other direction, given a dynamical system $(\mathbb{G},A,\alpha)$ which possesses Stone-von Neumann uniqueness, one may investigate what can be said of the elements of the dynamical system. We spend the remainder of this article pursuing these questions. As a preliminary step, we compute the spectrum of the bicrossed product.
Proposition 4.5.
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a $C^{*}$-algebra admitting a faithful state. Assume that there is a coaction $(\mathbb{G},A,\alpha)$ which is Stone-von Neumann unique. Then the spectrum of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ consists of a single point.
@@
We begin with a Stone-von Neumann Unique coaction $(\mathbb{G},A,\alpha)$, and must show that any two irreducible representations $\theta,\tilde{\theta}$ of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert spaces $H_{\theta},H_{\tilde{\theta}}$ are equivalent. First, consider $\theta\otimes 1$ as a representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert-$A$ module $H_{\theta}\otimes A$, similarly $\tilde{\theta}\otimes 1$. Using the canonical covariant representation of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\widehat{\alpha}^{\textnormal{op}})$
@@
$\begin{split}&\quad\;\bigl{\|}(\text{id}\otimes\Lambda_{\varpi})\bigl{(}\sum_{i}\xi_{i}\otimes a_{i}\bigr{)}\bigr{\|}^{2}=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\langle\Lambda_{\varpi}(a_{i})|\Lambda_{\varpi}(a_{j})\rangle=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\varpi(a_{i}^{*}a_{j})\\ &=\varpi\bigl{(}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{)}\leq\bigl{\|}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{\|}=\bigl{\|}\sum_{i}\xi_{i}\otimes a_{i}\bigr{\|}^{2}.\end{split}$
@@
$\begin{split}&\quad\;\bigl{\|}\tilde{U}\bigl{(}((\text{id}\otimes\Lambda_{\varpi})\Omega_{i})_{i\in S}\bigr{)}\bigr{\|}^{2}=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\bigr{)}_{i\in S}\bigr{\|}^{2}=\sum_{i\in S}\|(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\|^{2}\\ &=\sum_{i\in S}\varpi(\langle\zeta_{i},\zeta_{i}\rangle)=\varpi\bigl{(}\sum_{i\in S}\langle\zeta_{i},\zeta_{i}\rangle\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\zeta_{i})_{i\in S},(\zeta_{i})_{i\in S}\bigr{\rangle}\bigr{)}\\ &=\varpi\bigl{(}\bigl{\langle}U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)},U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)}\bigr{\rangle}\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\Omega_{i})_{i\in S},(\Omega_{i})_{i\in S}\bigr{\rangle}\bigr{)}=\sum_{i\in S}\varpi(\langle\Omega_{i},\Omega_{i}\rangle)\\ &=\sum_{i\in S}\langle(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}|(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\rangle=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\bigr{)}_{i\in S}\bigr{\|}^{2}\end{split}$
@@
The remaining conditions can be checked in a completely analogous way. This shows that $\bigoplus_{S}\theta\otimes 1$ and $\bigoplus_{S}\tilde{\theta}\otimes 1$ are equivalent. Consequently, there are multiples of $\theta$ and $\tilde{\theta}$ which are equivalent. But since $\theta,\tilde{\theta}$ are assumed to be irreducible, it follows that they are equivalent themselves ([DixmierC, Proposition 5.3.3]). In particular, they are equivalent as representations of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ and consequently this $C^{*}$-algebra has spectrum consisting of one point.
@@
With this in hand, we will now see that a complete characterization of coefficient algebras $A$ boils down to Naimark’s problem [naimark]. To this end, let us make a general observation that for the trivial coaction $(\mathbb{G},A,\operatorname{triv})$, we have
(7) $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\simeq(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A.$
@@
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a separable $C^{*}$-algebra, and assume that $(\mathbb{G},A,\operatorname{triv})$ enjoys Stone-von Neumann uniqueness. Then $A$ is an elementary $C^{*}$-algebra

### v2

@@ 3. Heisenberg Representations
Modular representations for group actions on operator algebras already appear in [HI] and [modular], emphasizing pointwise notions of covariant representations. Representations of actions lack such a pointwise notions, and require a distinct but parallel theory, which is more appropriate from the lens of locally compact quantum groups. Below, we offer a definition of Heisenberg representations which extends the acceptable dynamics and unifies the classical setting under the common theory of representations.
@@
Let $(\mathbb{G},A,\alpha)$ be a left dynamical system. A Heisenberg representation $(X,\rho^{c},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$ consists of a Hilbert $A$-module $X$ and non-degenerate *-homomorphisms
@@
be non-degenerate *-homomorphisms. The following are equivalent:
@@
    $(\widehat{\rho},\rho^{c})$ is a covariant representation of $(C^{u}_{0}(\widehat{\mathbb{G}}),\widehat{\mathbb{G}},({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u})$ on $X$.
@@
be non-degenerate *-homomorphisms. The following are equivalent:
@@
    $(\pi,\rho^{c})$ is a covariant representation of the right dynamical system $(A,\widehat{\mathbb{G}},{\rm triv})$ on $X$.
Here, ${\rm triv}\colon A\rightarrow M(A\otimes C_{0}(\widehat{\mathbb{G}}))$ denotes the right trivial action given by ${\rm triv}(a)=a\otimes 1$.
@@
Consider a Heisenberg $(\mathbb{G},A,{\rm triv})$-representation $(X,\rho^{c},\hat{\rho},\pi)$ comprising $A=\mathbb{C}$, $X=H$ a Hilbert space, and $\mathbb{G}=G$ a classical abelian locally compact group. Then $\rho^{c}$ and $\widehat{\rho}$ correspond to unitary representations $\widehat{U}$ and $U$ of $\widehat{G}$ and $G$ on $H$ respectively. The conditions (R1) and (R3) are automatically satisfied. Let ${\rm ev}_{g}\colon C_{0}(G)\to\mathbb{C}$ and ${\rm ev}_{\chi}\colon C_{0}(\widehat{G})\to\mathbb{C}$ be the evaluation functionals at points $g\in G$ and $\chi\in\widehat{G}$ respectively. After applying $({\rm ev}_{g}\otimes{\rm id}\otimes{\rm ev}_{\chi})$ to the defining equation of condition (R2), we obtain
@@
Rieffel refined and generalized Mackey’s pioneering work on induced representations of locally compact groups to the setting of of $C^{*}$-representation theory (see the introduction to [rinduced] for a nice history), which recovers the Mackey-Stone–von Neumann Theorem treated in Example 3.4 as a special case of Mackey’s Imprimitivity Theorem (follow the references in [Rieffel]*Paragraph Three). In the latter article, Rieffel defines Heisenberg $G$-modules for an arbitrary locally compact group $G$. These are exactly our Heisenberg $(\mathbb{G},A,\alpha)$-representations for $A=\mathbb{C}$ and $\mathbb{G}$ the dual to a classical locally compact group. The discerning will notice that the integrand in [Rieffel]*Equation (3) has striking similarities to the application of the left regular representation interpreted in the classical setting, an antecedent to our current work.
@@
In [mrw], the authors introduced the notion of a Heisenberg pair which Roy and Timmermann later generalized in [RT18]. Meyer, Roy, and Woronowicz define Heisenberg pairs in the context of quantum groups derived from manageable multiplicative unitaries (see e.g. [Wor96]); this concept is closely related to our Heisenberg representations in the sense of Kustermans and Vaes. Compare for instance the parallels between our Proposition 3.9 and the argument in [RT18, Theorem 6.2]. It is known that Woronowicz description of quantum groups captures the Kustermans and Vaes framework– that is, given a locally compact quantum group in the sense of this paper, one may construct a manageable multiplicative unitary of the type considered in [mrw].
We end this section with a result which relates our Heisenberg modular representations to dynamics of the crossed product. Much as full crossed products completely capture the dynamics of the associated system, our Heisenberg representations retain the data involved in crossed product constructions.
@@
Let $(\mathbb{G},A,\alpha)$ be a left dynamical system and $X$ a Hilbert $A$-module. There is a one-to-one correspondence between Heisenberg representations of $(\mathbb{G},A,\alpha)$ on $X$ and covariant representations of the left dynamical system $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$ on $X$.
@@
$=\chi((\widehat{\rho}\otimes{\rm id})({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u}(h))$
$=\chi((\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\widehat{\rho}(h)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*})$
@@
4. The Covariant Stone–von Neumann Theorem
@@
Let $(\mathbb{G},A,\alpha)$ be a left dynamical system. A Schrödinger representation is a quadruple $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha)$, where
@@
are the canonical surjective *-homomorphisms.
@@
$=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}(\mathrm{V}^{\widehat{\mathbb{G}}}_{23})^{*}\mathrm{V}^{\widehat{\mathbb{G}}}_{23}=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}$
$=(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}\,.$
@@
Using the machinery we have gathered so far, we can prove one of the main results of our work.
@@
Let $\left(\mathbb{G},A,\alpha\right)$ be a left dynamical system, where $A$ is an elementary $C^{*}$-algebra, $\mathbb{G}$ is a regular locally compact quantum group, and $\alpha$ is a maximal action. Then every Heisenberg $\left(\mathbb{G},A,\alpha\right)$-representation is a multiple of the Schrödinger $\left(\mathbb{G},A,\alpha\right)$-representation.
@@
Consider a Heisenberg $(\mathbb{G},A,\alpha)$-representation $(X,\rho^{\textnormal{c}},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$. Then Proposition 3.9 yields a covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ for $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$. In turn, the covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ determines a nondegenerate representation of the iterated crossed product
@@
Since $\alpha$ is maximal, we have an isomorphism $(\lambda^{c}\otimes 1)\ltimes(\hat{\lambda}\otimes 1)\ltimes\alpha\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\to\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$. Abbreviating the isomorphism by $\Phi$, we witness $X$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence with left action coming from the *-homomorphism $(\rho^{c}\ltimes\hat{\rho}\ltimes\pi)\circ\Phi^{-1}$.
Applying this analysis to the Schrödinger representation $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha),$ we can also see that the correspondence $L^{2}(\mathbb{G})\otimes A$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence. It is in fact an imprimitivity bimodule. To see this, notice that the map $\Phi$ in the previous paragraph agrees with the iterated crossed product *-homomorphism associated to the Schrödinger representation, so that the left action of $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ is given by the identity. Since $A$ and $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ are both simple, the fullness of $L^{2}(\mathbb{G})\otimes A$ is clear, and for the compatibility of the inner products, we compute for elementary tensors
@@
Thus we may apply Theorem 2.1 to deduce that $X$ is a multiple of $L^{2}(\mathbb{G})\otimes A$. We suppress the unitary which implements the equivalence $X\cong\bigoplus_{S}L^{2}(\mathbb{G})\otimes A$, and proceed by identifying these two modules.
@@
recovering the covariant representation $(\rho^{\textnormal{c}},\hat{\rho}\ltimes\pi)$ of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\hat{\alpha}^{\textnormal{op}}).$ Applying Proposition 3.9, we recover the Heisenberg $(\mathbb{G},A,\alpha)$-representation $(\oplus_{S}\big{(}L^{2}(\mathbb{G})\otimes A\big{)},\rho^{\textnormal{c}},\hat{\rho},\pi).$ Since each summand is invariant under the *-representations, this concludes the proof. ∎
Let us explicitly state the above result for $A=\mathbb{C}$ which can be thought of as a generalization of the Mackey–Stone–von Neumann theorem [Mackey] to quantum groups.
Corollary 4.4.
Let $\mathbb{G}$ be a regular locally compact quantum group and assume that the trivial of $\mathbb{G}$ on $\mathbb{C}$ is maximal. Then every Heisenberg $(\mathbb{G},\mathbb{C},{\rm triv})$-representation is a multiple of the Schrödinger $(\mathbb{G},\mathbb{C},{\rm triv})$-representation $(L^{2}(\mathbb{G}),\lambda^{c},\hat{\lambda},{\rm triv})$.
In light of Theorem 4.3, we introduce a new property for a left dynamical system $(\mathbb{G},A,\alpha)$, which is a quantum-dynamical analog of the behavior described in [Rieffel]*Theorem and [wcrossed]*Theorem 4.29.
@@
A dynamical system $(\mathbb{G},A,\alpha)$ is called Stone–von Neumann unique provided that every Heisenberg $(\mathbb{G},A,\alpha)$-representation is a multiple of the Schrödinger $(\mathbb{G},A,\alpha)$-representation.
By Theorem 4.3 if $\mathbb{G}$ is regular, $A$ is elementary and $\alpha$ is maximal, then $(\mathbb{G},A,\alpha)$ is Stone–von Neumann unique. In the other direction, given a dynamical system $(\mathbb{G},A,\alpha)$ which possesses Stone–von Neumann uniqueness, one may investigate what can be said of the elements of the dynamical system. We spend the remainder of this article pursuing these questions. The following key theorem is the main result of the paper, which relates the multiplicity result from Theorem 4.3 to the properties of the spectra and provides the essential tool for analyzing Stone-von Neumann unique dynamical systems.
Theorem 4.6.
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a $C^{*}$-algebra admitting a faithful state. Assume that there is a dynamical $(\mathbb{G},A,\alpha)$ which is Stone–von Neumann unique. Then the spectrum of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ consists of a single point.
@@
We begin with a Stone–von Neumann unique dynamical system $(\mathbb{G},A,\alpha)$, and must show that any two irreducible representations $\theta,\tilde{\theta}$ of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert spaces $H_{\theta},H_{\tilde{\theta}}$ are equivalent. First, consider $\theta\otimes 1$ as a representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert-$A$ module $H_{\theta}\otimes A$, similarly $\tilde{\theta}\otimes 1$. Using the canonical covariant representation of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\widehat{\alpha}^{\textnormal{op}})$
@@
$\begin{split}\bigl{\|}(\text{id}\otimes\Lambda_{\varpi})\bigl{(}\sum_{i}\xi_{i}\otimes a_{i}\bigr{)}\bigr{\|}^{2}&=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\langle\Lambda_{\varpi}(a_{i})|\Lambda_{\varpi}(a_{j})\rangle=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\varpi(a_{i}^{*}a_{j})\\ &=\varpi\bigl{(}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{)}\leq\bigl{\|}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{\|}\\ &=\bigl{\|}\sum_{i}\xi_{i}\otimes a_{i}\bigr{\|}^{2}.\end{split}$
@@
$\begin{split}\bigl{\|}\tilde{U}\bigl{(}((\text{id}\otimes\Lambda_{\varpi})\Omega_{i})_{i\in S}\bigr{)}\bigr{\|}^{2}&=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\bigr{)}_{i\in S}\bigr{\|}^{2}=\sum_{i\in S}\|(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\|^{2}\\ &=\sum_{i\in S}\varpi(\langle\zeta_{i},\zeta_{i}\rangle)=\varpi\bigl{(}\sum_{i\in S}\langle\zeta_{i},\zeta_{i}\rangle\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\zeta_{i})_{i\in S},(\zeta_{i})_{i\in S}\bigr{\rangle}\bigr{)}\\ &=\varpi\bigl{(}\bigl{\langle}U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)},U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)}\bigr{\rangle}\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\Omega_{i})_{i\in S},(\Omega_{i})_{i\in S}\bigr{\rangle}\bigr{)}\\ &=\sum_{i\in S}\varpi(\langle\Omega_{i},\Omega_{i}\rangle)=\sum_{i\in S}\langle(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}|(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\rangle\\ &=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\bigr{)}_{i\in S}\bigr{\|}^{2}.\end{split}$
@@
The remaining conditions can be checked in a completely analogous way. This shows that $\bigoplus_{S}\theta\otimes 1$ and $\bigoplus_{S}\tilde{\theta}\otimes 1$ are equivalent. Consequently, they are multiples of $\theta$ and $\tilde{\theta}$ which are equivalent. But since $\theta,\tilde{\theta}$ are assumed to be irreducible, it follows that they are equivalent themselves ([DixmierC, Proposition 5.3.3]). In particular, they are equivalent as representations of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ and consequently this $C^{*}$-algebra has spectrum consisting of one point. ∎
@@
With this in hand, we will now see that a complete characterization of coefficient algebras $A$ boils down to Naimark’s problem [naimark]. To this end, let us make a general observation that for the trivial action $(\mathbb{G},A,\operatorname{triv})$, we have
(10) $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\cong(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A.$
@@
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a separable $C^{*}$-algebra, and assume that $(\mathbb{G},A,\operatorname{triv})$ enjoys Stone–von Neumann uniqueness. Then $A$ is an elementary $C^{*}$-algebra.